the morena family drove to yoesmite national park in 3 hours 24 minutes. on the way home, they averaged5 miles per hour less and the trip took 3 hours 42 min. how far from yosemite national park do they live?
Answer by richwmiller(17219)

24/60==4/10=.4
42/60=7/10=.7
3.4x=3.7*(x-5)
x=61 2/3
x+5=66 2/3
3.4*61 2/3
209 2/3 miles